sqrt (x/3) sqrt (5/x-2) Multiply.
Answer by profemmanuel2q@yahoo.com(15)

sqrt (x/3) sqrt (5/x-2) Multiply.
This question too is easy but you need to think
square of (x/3) is (x^2/9)
and the square of (5/x-2) is 25/ (x - 2)^2
So if you want to multiply this, you will get:
(x^2/9) X 25/(x - 2)^2
x^2 X 25
9 (x – 2)^2
So we have 25x^2 or u can expand and multiply 9(x – 2)^2
9(x – 2)^2

9(x – 2)^2 = 9 (x^2 – 4x + 4) = 9x^2 – 36x + 36
So the final answer is 25x^2 or 25x^2
9(x – 2)^2 (9x^2 – 36x + 36)